OpenStudy (anonymous):

If h(x) = −3x − 10, find h(−3).

OpenStudy (anonymous):

h(-3)=-1

hero (hero):

-3(-3) - 10 = 9 - 10 = -1

OpenStudy (anonymous):

h(-3) = -3 (-3) -10 = 9 - 10 so = -1

hero (hero):